# Efficient Neuro-Symbolic Retrieval-Augmented Generation through Adaptive Query Routing

### v1

@@ Abstract
Keywords:
Neuro-symbolic reasoning Adaptive routing Retrieval-Augmented Generation Hybrid AI Resource-aware computation LLM efficiency
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems address factual inconsistencies in Large Language Models by grounding generation in external knowledge, yet they face a fundamental efficiency problem: simple queries consume computational resources equivalent to complex multi-hop reasoning tasks [4, 12]. Current RAG architectures apply uniform processing pipelines regardless of query complexity, leading to substantial computational overhead from dense neural retrieval and quadratic attention scaling [2]. This one-size-fits-all approach becomes increasingly problematic as model scale and deployment demands grow, particularly given rising concerns about AI’s energy footprint and the need for sustainable deployment.
@@
  • Cross-architecture validation: Empirical validation across Llama-3.2-3B [6] and Mistral-7B [7], demonstrating that disabling adaptive logic causes 169-1151% processing time increases.
@@ 2 Related Work
Retrieval-Augmented Generation enriches LLMs with external knowledge [4, 12] but often requires trade-offs between high computational cost (neural-only) and brittle logic (symbolic-only). Hybrid RAG methods attempt to bridge this gap through various integration strategies.
Several hybrid approaches incorporate symbolic knowledge but lack dynamic adaptation to query complexity or system resources. RuleRAG [1] and RuAG [11] inject handcrafted or learned rules into retrieval and generation pipelines, while HybridRAG [8] statically fuses knowledge graphs with vector contexts without considering query-specific requirements or computational constraints.
Another line of work focuses on graph-based reasoning approaches. Graph-based methods like CRP-RAG [9] construct reasoning graphs for multi-hop QA but rely on multiple sequential LLM calls and fixed structures. CDF-RAG [5] adds causal graphs with reinforcement learning-driven query refinement, but its RL loop introduces approximately 100ms latency per query without optimizing overall system throughput.
SymRAG differs fundamentally by implementing real-time adaptive routing based on complexity assessment and resource monitoring. Unlike static integration approaches, SymRAG dynamically selects optimal processing pathways (symbolic, neural, or hybrid) for each query, achieving high accuracy (97.6-100% exact match) with efficient resource utilization explicitly reported across multiple metrics. A detailed comparison of RAG frameworks emphasizing SymRAG’s unique contributions is provided in Appendix 0.A.11.
@@ 3 SymRAG: Adaptive Query Routing Framework
Architectural Overview: SymRAG integrates four core components—Graph-Based Symbolic Reasoner, Neural Retriever, Hybrid Integrator, and Resource-Aware System Control Manager—operating across Query Processing, Hybrid Processing, and Resource Management stages. The system employs utility-based path selection that balances accuracy, latency, and computational costs through continuous learning and threshold adaptation. Unlike static integration approaches, SymRAG exhibits meta-learning characteristics where optimal resource allocation patterns emerge through experience, enabling both superior performance and computational sustainability. Complete architectural details, component interactions, and system implementation specifics are provided in Appendix 0.A.2.
@@
Definition 1 (Query Complexity $\kappa(q)$)
@@
Definition 2 (Path Selection Policy)
@@ 4 Experimental Setup
We selected datasets testing fundamentally different reasoning capabilities to assess SymRAG’s versatility and adaptive behavior. HotpotQA [10] provides a challenging multi-hop question answering benchmark requiring reasoning across multiple documents, with 1,000 development set queries demanding integration of information from multiple supporting facts while testing complex inference chains and handling textual ambiguity. DROP [3] offers a discrete reasoning benchmark evaluating numerical operations, comparisons, and span extraction over paragraphs, with 1,000 queries providing robust statistical power for discrete mathematical operations requiring precise understanding and manipulation of quantities, dates, and entities.
@@
Primary baselines include Neural-Only RAG using solely the NeuralRetriever component representing current state-of-the-art approaches, and Symbolic-Only processing exclusively through the GraphSymbolicReasoner demonstrating pure rule-based reasoning capabilities. Ablation configurations systematically remove or modify key components: disabling adaptive logic forces all queries through hybrid paths regardless of complexity or resource state, removing few-shot prompting assesses impact on discrete reasoning tasks, and comparing static versus dynamic rules evaluates automated rule learning benefits. Complete experimental setup details including hardware specifications and hyperparameters are provided in Appendix 0.A.10.
@@ 5.2 Cross-Model Generalizability and Adaptive Logic Impact
To validate SymRAG’s generalizability and the critical importance of adaptive logic, we conducted comprehensive ablation studies across both LLM architectures. Table 2 presents the dramatic impact of disabling adaptive logic, establishing universal efficiency criticality despite model-specific variations (see Appendix 0.A.7 for a visual representation).
Table 2: Cross-Model Impact of Disabling Adaptive Logic
Model Configuration EM (%) F1 (%) Avg Time (s) Acc. Impact Speed Impact p-value Llama-3.2-3B With Adaptive 89.2 89.4 0.985 No Adaptive 75.7 75.9 2.645 –15.1 % +168.6 % $<$0.001 Effect Size (d) 0.78 0.78 Large Large Mistral-7B With Adaptive 89.2 89.4 2.443 No Adaptive 86.2 86.1 25.847 –3.4 % +958.0 % $<$0.001 Effect Size (d) 0.03 0.03 –0.60 Small Very Large
@@ 5.3 Comparative Performance Analysis
SymRAG achieves competitive accuracy while maintaining superior resource efficiency through adaptive path selection. The system’s ability to dynamically route queries based on complexity enables both performance consistency and energy optimization compared to uniform neural processing approaches. Statistical significance testing and detailed comparative analysis across different model architectures are provided in Appendix 0.A.9.
@@ 5.4 Component Contribution Analysis
Resource usage patterns show interesting temporal dynamics, with CPU usage actually decreasing over time during extended processing sessions (0.9% reduction for HotpotQA, 2.68% for DROP over 1,000 queries; see Appendix 0.A.3.1 for detailed trends). This improvement occurs through increasingly effective caching mechanisms, adaptive threshold stabilization around optimal values, and more efficient resource allocation as the system learns typical requirements for different query types. Processing paths show distinct resource patterns, with hybrid achieving better efficiency than pure neural approaches. Full component analysis results are provided in Appendix 0.A.8.1.
@@ 6 Discussion
Extended future directions including multi-modal reasoning and theoretical foundations are detailed in Appendix 0.A.13.
@@ 7 Conclusion
Appendix 0.A Appendix: Supplementary Material
@@
0.A.1 Mathematical Foundations & Algorithms
@@
0.A.1.1 Query Complexity Assessment
The query complexity score $\kappa(q)$ introduced in Definition 1 integrates multiple linguistic and structural indicators through a principled mathematical framework. The attention values $A(q)$ are computed as the mean of final layer attention head values from prajjwal1/bert-tiny when processing query $q$:
@@ 0.A.1.1 Query Complexity Assessment
0.A.1.2 Resource State Dynamics and System Load Modeling
@@
Definition 3 (Resource State Vector and Pressure)
@@
0.A.1.3 Path Utility and Optimization Framework
@@
Definition 4 (Path Utility Function)
@@
0.A.1.4 Fusion Confidence and Integration Assessment
@@
Definition 5 (Fusion Confidence Assessment)
@@
$C_{fusion}=\begin{cases}\min(C_{sym},\,C_{neur})\cdot\beta_{agree},&\begin{aligned} &\text{if }\;\text{TypeMatch}(A_{sym},\,A_{neur})=1\\ &\quad\wedge\;\text{ValueMatch}(A_{sym},\,A_{neur})=1\end{aligned}\\ \max(C_{sym},\,C_{neur})\cdot\beta_{conflict},&\begin{aligned} &\text{if }\;\text{TypeMatch}(A_{sym},\,A_{neur})=1\\ &\quad\wedge\;\text{ValueMatch}(A_{sym},\,A_{neur})=0\end{aligned}\\ \frac{C_{sym}+C_{neur}}{2}\cdot\beta_{mismatch},&\text{if }\;\text{TypeMatch}(A_{sym},\,A_{neur})=0\end{cases}$
@@
0.A.1.5 Adaptive Path Selection
We use dynamic thresholds to choose between symbolic ($P_{S}$), neural ($P_{N}$), or hybrid ($P_{H}$) paths depending on query complexity and system resource pressure. The default path is $P_{H}$. These thresholds ($T_{low\_\kappa}$, $T_{high\_\kappa}$ for complexity; $T_{low\_R}$, $T_{high\_R}$ for resource pressure) are themselves subject to optimization (see Algorithm 2). For example, if both effective query complexity $\kappa_{eff}(q)$ and resource pressure $R_{p}(t)$ are below their respective low thresholds, a symbolic path is favored. Conversely, if either complexity or pressure meets or exceeds their high thresholds, the neural path is generally chosen. The detailed logic is presented in Algorithm 1.
Algorithm 1 Adaptive Path Selection Logic
Input: Query complexity $\kappa_{eff}(q)$, resource pressure $R_{p}(t)$, dataset type $D_{t}\in\{\text{DROP},\text{HotpotQA}\}$
@@ 1. Set $P^{*}\leftarrow P_{N}$
Algorithm 2 Adaptive Threshold Optimization Logic
Input: Current thresholds $T_{\text{curr}}$ (e.g., a structure with fields like $T_{\text{curr}}.T_{\text{low}\_\kappa}$), resource pressure $R_{p}(t)$, path performance stats $S_{\text{paths}}$
Output: Updated thresholds $T_{\text{updated}}$
@@
    Initialize $T_{\text{updated}}\leftarrow T_{\text{curr}}$, $\delta\leftarrow 0.05$
@@
      Set $T_{\text{updated}}.T_{\text{low}\_\kappa}\leftarrow\min(0.6,T_{\text{updated}}.T_{\text{low}\_\kappa}+\delta)$
@@
      Set $T_{\text{updated}}.T_{\text{high}\_\kappa}\leftarrow\max(0.6,T_{\text{updated}}.T_{\text{high}\_\kappa}-\delta)$
@@
      Set $T_{\text{updated}}.T_{\text{low}\_\kappa}\leftarrow\max(0.2,T_{\text{updated}}.T_{\text{low}\_\kappa}-\delta)$
@@
      Set $T_{\text{updated}}.T_{\text{high}\_\kappa}\leftarrow\min(0.9,T_{\text{updated}}.T_{\text{high}\_\kappa}+\delta)$
@@
        If $P=P_{N}$ and $T_{\text{updated}}.T_{\text{high}\_\kappa}>0.6$: Set $T_{\text{updated}}.T_{\text{high}\_\kappa}\leftarrow\max(0.6,T_{\text{updated}}.T_{\text{high}\_\kappa}-\delta)$
@@
        If $P=P_{S}$ and $T_{\text{updated}}.T_{\text{low}\_\kappa}>0.2$: Set $T_{\text{updated}}.T_{\text{low}\_\kappa}\leftarrow\max(0.2,T_{\text{updated}}.T_{\text{low}\_\kappa}-\delta)$
@@
    Return $T_{\text{updated}}$
0.A.1.6 Dynamic Threshold Optimization
The threshold adaptation mechanism enables SymRAG to learn optimal routing strategies through experience, operating through resource-based and performance-based feedback loops (see Algorithm 2).
@@ 0.A.1.6 Dynamic Threshold Optimization
0.A.2 Detailed System Architecture and Component Integration
@@
0.A.3 Architecture Components & Resource Management
0.A.3.1 Resource State Vector Formulation
@@
$R(t)=\begin{bmatrix}\text{CPU}(t)\\ \text{GPU}(t)\\ \text{MEM}(t)\\ \text{Power}(t)\end{bmatrix}\in[0,1]^{4}$
@@
0.A.3.2 Symbolic-Neural Integration Framework
@@
0.A.4 System Analysis & Dynamic Rule Extraction
@@ 0.A.5 Fusion Mechanisms
0.A.6 Experimental Validation & Component Analysis
@@
SymRAG achieves 1.6-2.4 percentage point improvements over Neural-Only baselines, representing hundreds of additional correct answers in production scenarios. The catastrophic failure of Symbolic-Only approaches (31.6% for DROP, 42.1% for HotpotQA) validates the necessity of neural components, while symbolic reasoning’s exceptional speed (0.442s for DROP, 0.089s for HotpotQA) highlights its value when integrated intelligently. All performance differences achieve statistical significance (p < 0.05) with effect sizes ranging from medium (d = 0.34) to large (d = 0.89).
0.A.7 Ablation Studies & Adaptive Logic Criticality

### v2

@@ Abstract
Accepted at 19th International Conference on Neurosymbolic Learning and Reasoning (NeSy 2025)
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems address factual inconsistencies in Large Language Models by grounding generation in external knowledge, yet they face a fundamental efficiency problem: simple queries consume computational resources equivalent to complex multi-hop reasoning tasks (Fan et al., 2024; Zhao et al., 2024). Current RAG architectures apply uniform processing pipelines regardless of query complexity, leading to substantial computational overhead from dense neural retrieval and quadratic attention scaling (Chitty-Venkata et al., 2024). This one-size-fits-all approach becomes increasingly problematic as model scale and deployment demands grow, particularly given rising concerns about AI’s energy footprint and the need for sustainable deployment.
@@
  • Cross-architecture validation: Empirical validation across Llama-3.2-3B (Llama, 2024) and Mistral-7B (MistralAI, 2024), demonstrating that disabling adaptive logic causes 169-1151% processing time increases.
@@ 2 Related Work
Retrieval-Augmented Generation enriches LLMs with external knowledge (Fan et al., 2024; Zhao et al., 2024) but often requires trade-offs between high computational cost (neural-only) and brittle logic (symbolic-only). Hybrid RAG methods attempt to bridge this gap through various integration strategies.
Several hybrid approaches incorporate symbolic knowledge but lack dynamic adaptation to query complexity or system resources. RuleRAG (Chen et al., 2024) and RuAG (Zhang et al., 2024) inject handcrafted or learned rules into retrieval and generation pipelines, while HybridRAG (Sarmah et al., 2024) statically fuses knowledge graphs with vector contexts without considering query-specific requirements or computational constraints.
Another line of work focuses on graph-based reasoning approaches. Graph-based methods like CRP-RAG (Xu et al., 2024) construct reasoning graphs for multi-hop QA but rely on multiple sequential LLM calls and fixed structures. CDF-RAG (Khatibi et al., 2025) adds causal graphs with reinforcement learning-driven query refinement, but its RL loop introduces approximately 100ms latency per query without optimizing overall system throughput.
SymRAG differs fundamentally by implementing real-time adaptive routing based on complexity assessment and resource monitoring. Unlike static integration approaches, SymRAG dynamically selects optimal processing pathways (symbolic, neural, or hybrid) for each query, achieving high accuracy (97.6-100% exact match) with efficient resource utilization explicitly reported across multiple metrics. A detailed comparison of RAG frameworks emphasizing SymRAG’s unique contributions is provided in Appendix A.11.
@@ 3 SymRAG: Adaptive Query Routing Framework
Architectural Overview: SymRAG integrates four core components—Graph-Based Symbolic Reasoner, Neural Retriever, Hybrid Integrator, and Resource-Aware System Control Manager—operating across Query Processing, Hybrid Processing, and Resource Management stages. The system employs utility-based path selection that balances accuracy, latency, and computational costs through continuous learning and threshold adaptation. Unlike static integration approaches, SymRAG exhibits meta-learning characteristics where optimal resource allocation patterns emerge through experience, enabling both superior performance and computational sustainability. Complete architectural details, component interactions, and system implementation specifics are provided in Appendix A.2.
@@
Definition 3.1 (Query Complexity $\kappa(q)$).
@@
Definition 3.2 (Path Selection Policy).
@@ 4 Experimental Setup
We selected datasets testing fundamentally different reasoning capabilities to assess SymRAG’s versatility and adaptive behavior. HotpotQA (Yang et al., 2018) provides a challenging multi-hop question answering benchmark requiring reasoning across multiple documents, with 1,000 development set queries demanding integration of information from multiple supporting facts while testing complex inference chains and handling textual ambiguity. DROP (Dua et al., 2019) offers a discrete reasoning benchmark evaluating numerical operations, comparisons, and span extraction over paragraphs, with 1,000 queries providing robust statistical power for discrete mathematical operations requiring precise understanding and manipulation of quantities, dates, and entities.
@@
Primary baselines include Neural-Only RAG using solely the NeuralRetriever component representing current state-of-the-art approaches, and Symbolic-Only processing exclusively through the GraphSymbolicReasoner demonstrating pure rule-based reasoning capabilities. Ablation configurations systematically remove or modify key components: disabling adaptive logic forces all queries through hybrid paths regardless of complexity or resource state, removing few-shot prompting assesses impact on discrete reasoning tasks, and comparing static versus dynamic rules evaluates automated rule learning benefits. Complete experimental setup details including hardware specifications and hyperparameters are provided in Appendix A.10.
@@ 5.2 Cross-Model Generalizability and Adaptive Logic Impact
To validate SymRAG’s generalizability and the critical importance of adaptive logic, we conducted comprehensive ablation studies across both LLM architectures. Table 2 presents the dramatic impact of disabling adaptive logic, establishing universal efficiency criticality despite model-specific variations (see Appendix A.7 for a visual representation).
Table 2: Cross‐Model Impact of Disabling Adaptive Logic
Model Configuration EM (%) F1 (%) Avg Time (s) Acc. Impact Speed Impact p‐value Llama-3.2-3B With Adaptive 89.2 89.4 0.985 No Adaptive 75.7 75.9 2.645 –15.1 % +168.6 % $<$0.001 Effect Size (d) 0.78 0.78 Large Large Mistral-7B With Adaptive 89.2 89.4 2.443 No Adaptive 86.2 86.1 25.847 –3.4 % +958.0 % $<$0.001 Effect Size (d) 0.03 0.03 –0.60 Small Very Large
@@ 5.3 Comparative Performance Analysis
SymRAG achieves competitive accuracy while maintaining superior resource efficiency through adaptive path selection. The system’s ability to dynamically route queries based on complexity enables both performance consistency and energy optimization compared to uniform neural processing approaches. Statistical significance testing and detailed comparative analysis across different model architectures are provided in Appendix A.9.
@@ 5.4 Component Contribution Analysis
Resource usage patterns show interesting temporal dynamics, with CPU usage actually decreasing over time during extended processing sessions (0.9% reduction for HotpotQA, 2.68% for DROP over 1,000 queries; see Appendix A.3.1 for detailed trends). This improvement occurs through increasingly effective caching mechanisms, adaptive threshold stabilization around optimal values, and more efficient resource allocation as the system learns typical requirements for different query types. Processing paths show distinct resource patterns, with hybrid achieving better efficiency than pure neural approaches. Full component analysis results are provided in Appendix A.8.1.
@@ 6 Discussion
Extended future directions including multi-modal reasoning and theoretical foundations are detailed in Appendix A.13.
@@ 7 Conclusion
Appendix A Appendix: Supplementary Material
@@
A.1 Mathematical Foundations & Algorithms
@@
A.1.1 Query Complexity Assessment
The query complexity score $\kappa(q)$ introduced in Definition 3.1 integrates multiple linguistic and structural indicators through a principled mathematical framework. The attention values $A(q)$ are computed as the mean of final layer attention head values from prajjwal1/bert-tiny when processing query $q$:
@@ A.1.1 Query Complexity Assessment
A.1.2 Resource State Dynamics and System Load Modeling
@@
Definition A.1 (Resource State Vector and Pressure).
@@
A.1.3 Path Utility and Optimization Framework
@@
Definition A.2 (Path Utility Function).
@@
A.1.4 Fusion Confidence and Integration Assessment
@@
Definition A.3 (Fusion Confidence Assessment).
@@
[…]
@@
A.1.5 Adaptive Path Selection
We use dynamic thresholds to choose between symbolic ($P_{S}$), neural ($P_{N}$), or hybrid ($P_{H}$) paths depending on query complexity and system resource pressure. The default path is $P_{H}$. These thresholds ($T_{low\_\kappa}$, $T_{high\_\kappa}$ for complexity; $T_{low\_R}$, $T_{high\_R}$ for resource pressure) are themselves subject to optimization (see Algorithm LABEL:alg:threshold_optimization). For example, if both effective query complexity $\kappa_{eff}(q)$ and resource pressure $R_{p}(t)$ are below their respective low thresholds, a symbolic path is favored. Conversely, if either complexity or pressure meets or exceeds their high thresholds, the neural path is generally chosen. The detailed logic is presented in Algorithm LABEL:alg:path-selection.
[…]
alg:path-selection Adaptive Path Selection Logic Input: Query complexity $\kappa_{eff}(q)$, resource pressure $R_{p}(t)$, dataset type $D_{t}\in\{\text{DROP},\text{HotpotQA}\}$
@@ 1. Set $P^{*}\leftarrow P_{N}$
[…]
alg:threshold_optimization Adaptive Threshold Optimization Logic Input: Current thresholds $T_{curr}$ (e.g., a structure with fields like $T_{curr}.T_{low\_\kappa}$), resource pressure $R_{p}(t)$, path performance stats $S_{paths}$
Output: Updated thresholds $T_{updated}$
@@
    Initialize $T_{updated}\leftarrow T_{curr}$, $\delta\leftarrow 0.05$
@@
      Set $T_{updated}.T_{low\_\kappa}\leftarrow\min(0.6,T_{updated}.T_{low\_\kappa}+\delta)$
@@
      Set $T_{updated}.T_{high\_\kappa}\leftarrow\max(0.6,T_{updated}.T_{high\_\kappa}-\delta)$
@@
      Set $T_{updated}.T_{low\_\kappa}\leftarrow\max(0.2,T_{updated}.T_{low\_\kappa}-\delta)$
@@
      Set $T_{updated}.T_{high\_\kappa}\leftarrow\min(0.9,T_{updated}.T_{high\_\kappa}+\delta)$
@@
        If $P=P_{N}$ and $T_{updated}.T_{high\_\kappa}>0.6$: Set $T_{updated}.T_{high\_\kappa}\leftarrow\max(0.6,T_{updated}.T_{high\_\kappa}-\delta)$
@@
        If $P=P_{S}$ and $T_{updated}.T_{low\_\kappa}>0.2$: Set $T_{updated}.T_{low\_\kappa}\leftarrow\max(0.2,T_{updated}.T_{low\_\kappa}-\delta)$
@@
    Return $T_{updated}$
A.1.6 Dynamic Threshold Optimization
The threshold adaptation mechanism enables SymRAG to learn optimal routing strategies through experience, operating through resource-based and performance-based feedback loops (see Algorithm LABEL:alg:threshold_optimization).
@@ A.1.6 Dynamic Threshold Optimization
A.2 Detailed System Architecture and Component Integration
@@
A.3 Architecture Components & Resource Management
A.3.1 Resource State Vector Formulation
@@
[…]
@@
A.3.2 Symbolic-Neural Integration Framework
@@
A.4 System Analysis & Dynamic Rule Extraction
@@ A.5 Fusion Mechanisms
A.6 Experimental Validation & Component Analysis
@@
SymRAG achieves 1.6-2.4 percentage point improvements over Neural-Only baselines, representing hundreds of additional correct answers in production scenarios. The catastrophic failure of Symbolic-Only approaches (31.6% for DROP, 42.1% for HotpotQA) validates the necessity of neural components, while symbolic reasoning’s exceptional speed (0.442s for DROP, 0.089s for HotpotQA) highlights its value when integrated intelligently. All performance differences achieve statistical significance (p ¡ 0.05) with effect sizes ranging from medium (d = 0.34) to large (d = 0.89).
A.7 Ablation Studies & Adaptive Logic Criticality